The 59.4-g sample of the compound X4O6 contains 14.4 g of oxygen atoms. What is the molar mass of element X?
1
Expert's answer
2016-11-08T14:56:08-0500
The sample of the compound contains 14.4 g of oxygen than it contains (59.4-14.4)=45.0 g of element X. 1 mole of the compound contains (16.0*6)=96.0 g of oxygen. Therefore 1 mole of the compound contains (96.0*45.0)/14.4 = 300.0 g of element X. As 1 mole of the compound contains 4 moles of atoms of X, than molar mass of element X is 300.0/4 = 75.0 g/mol. We can in the periodic table that element X is Arsenic (As) and our compound is As4O6.

Answer:
The molar mass of element X is 75 g/mol. It is Arsenic.